OpenStudy (anonymous):

What is the surface area of a circle that has a diameter of 14 inches?

OpenStudy (anonymous):

http://math.about.com/library/blmeasurement.htm

OpenStudy (anonymous):

u could use the fomula which r^2*pi

OpenStudy (anonymous):

S=22/7*7*7=154sq.inch

OpenStudy (anonymous):

44 inches

OpenStudy (anonymous):

if the surface area is what is asked then it should be pi*r^2 and not 2pi*r

Directrix (directrix):

A circle does not have surface area. A sphere has surface area.

Directrix (directrix):

@erojda If you want to find the area of a circle, the A = pi * r ^2 = pi* 7^2 = 49 pi square inches. If you want to find the surface area of a sphere, SA = 4 * pi* r^2 = 4 * pi * 7^2 = 196 pi square inches. Note that radius = 1/2 of diameter. r = 7 if d = 14.